if a simple harmonic oscillator has got a displacement of 0.02m and acceleration equal to 2.0ms at any time the anguular frequency of the oscillator is

Grade:11

1 Answers

Adi
6 years ago
X= 0.02 ma= 2 m/s^(2)Now, a= xw^(2) ...(magnitude) Then, w = (a/x)^(1/2) = (2/0.02)^(1/2) = (100)^(1/2) = 10 rad/s